OpenStudy (anonymous):

the sum of two consecutive integers is 129 what is the value of the second integer?

OpenStudy (anonymous):

n-1 + n = 129

OpenStudy (anonymous):

n+(n+1)=129

OpenStudy (saifoo.khan):

suppose that number be x the next number be x + 1 finally, x + (x+1) = 129 2x = 128 x = 128/2 x = Answer.

OpenStudy (anonymous):

It's asking for the value of the second integer, not the first.

OpenStudy (anonymous):

so add one.. it's going to be okay :)

OpenStudy (saifoo.khan):

let her solve too!

OpenStudy (saifoo.khan):

Answer + 1 will give u the 2nd interger.

OpenStudy (saifoo.khan):

integer*

OpenStudy (anonymous):

so n is 64?

OpenStudy (saifoo.khan):

64 is the first number.. whts the second?

OpenStudy (anonymous):

65?

OpenStudy (saifoo.khan):

Right.

OpenStudy (anonymous):

yes

OpenStudy (saifoo.khan):

64+65 = 129

OpenStudy (anonymous):

yay ok thnx(:

OpenStudy (saifoo.khan):

welkx/